OpenStudy (anonymous):

The cost of a drive-in movie is $y per vehicle. A group of friends in a van shared the admission cost by paying $0.40 each. If 6 more friends had gone along, everyone would have paid only $0.25 each. What is the value of $y? (a) $4 (b) $6 (c) $8 (d) $10 (e) $12

OpenStudy (anonymous):

0.4x = y....(i) 0.25(x+6) =y.......(ii) equate equation (i) and (ii) 0.4x = 0.25(x+6) 0.4x = 0.25x+1.5 0.4x-0.25x = 1.5 0.15x = 1.5 x = 1.5/0.15 x = 10 y = 0.4x y=??

OpenStudy (anonymous):

$4?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you welcome